OpenStudy (anonymous):

A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 1.5 meters and a height of 1.25 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste? 14 13 5 4

OpenStudy (anonymous):

Can you solve \( n \times \frac 9 {16} \times \pi \times 1.25 \ge 10 \)?

OpenStudy (anonymous):

where are you getting 9/16 from

OpenStudy (anonymous):

\( 1.5= \frac 3 2\); Radius \(^2 = \frac 9 {16} \)

OpenStudy (anonymous):

i got 2.2. so now do I plug this in

OpenStudy (anonymous):

Yes, that's right!

OpenStudy (anonymous):

when i put it all in I get 4.8 or I just rounded it to 5

OpenStudy (anonymous):

you got that right :)

OpenStudy (anonymous):

thansk! hope i get a good grade on this

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

Yep. :)